OpenStudy (ilovecake):

Tessa bought stock in a restaurant for $233.00. Her stock is now worth $321.54. What is the percentage increase of the value of Tessa's stock? 89% 72% 38% 28%

OpenStudy (campbell_st):

use Change in price -------------- x 100 original price

OpenStudy (ilovecake):

Medal and fan.

OpenStudy (ilovecake):

@campbell_st what

OpenStudy (campbell_st):

change in price = 321.54 - 233.00 original price = 233 so (change in price)/(original price) * 100 =

OpenStudy (ilovecake):

I still don't get it....

OpenStudy (ilovecake):

Thanks for no help.

OpenStudy (campbell_st):

ok... by what amount of money did the share price go up...?

OpenStudy (campbell_st):

@Ilovecake that's ok... I'm here the help your understanding... and not provide you with answers... you need to do some work as well... change in price = 321.54 - 233.00 that's the start then divide the answer by the original price $233 you will have a decimal, multiply it by 100 to get a percentage.

OpenStudy (ilovecake):

some else gave me the answer.